) The Master (Static) budget for Jack Daniels Inc. is presented below. For June, Jack Daniels...

  1. ) The Master (Static) budget for Jack Daniels Inc. is presented below. For June, Jack Daniels Inc. manufactured and sold 920 units for $835. During this month the company incurred $460,000 total variable expenses and $180,000 total fixed expenses.
    1. Required
      1. Prepare a flexible budget performance report for June including all variances.

         Master (Static) Budget

Units

1,000

Sales

$800,000

Variable costs

450,000

Contribution margin

$350,000

Fixed costs

150,000

Operating income

$200,000

  1. Jack Daniels Inc. used 3,450 pounds of aluminum in June to manufacture 920 units. The company paid $28.50 per pound during the month to purchase aluminum. On June 1, the company had 50 pounds of aluminum on hand. At the end of June, the company had only 30 pounds of aluminum in its warehouse. Daniels used 4,340 direct labor hours in June, at an average cost of $41.50 per hour.  The standard for aluminum is 4 pounds per unit at a standard cost of $25.00 per pound. The standard for direct labor is 5 hours per unit at a standard rate of $40 per hour.
    1. Compute for June, Jack Daniels Inc.
      1. Purchase price variance for aluminum
      2. Direct labor rate and efficiency variance
    2. Daniels deployed new strategies for both aluminum and labor. For aluminum and labor they invested in better inputs. Did management make good decisions?
Answer #1

a.

Flexible Budget Performance Report

For The Month Of June

Actual results Revenue and spending variance Flexible budget Activity variance Master budget
Units 920 920 1,000
Sales $768,200 $32,200 F $736,000 (920*$800,000/1,000) $64,000 U $800,000
Variable cost 460,000 46,000 U 414,000 (920*$450,000/1,000) 36,000 F 450,000
Contribution margin 308,200 13,800 U 322,000 28,000 F 350,000
Fixed costs 180,000 30,000 U 150,000 0 150,000
Operating income $128,200 $43,800 U $172,000 $28,000 U $200,000

b.

i. Purchase price variance = Actual quantity*Standard price - Actual quantity*Actual price

Actual quantity purchased = 3450+30-50 = 3430 units

Purchase price variance = 3,430*$25 - 3,430*$28.50

Purchase price variance = $85,750 - 97,755 = $12,005 Unfavorable

Material usage variance = Standard quantity*Standard price - Actual quantity*Standard price

Material usage variance = 920*4*$25 - 3,450*$25

Material usage variance = $92,000 - 86,250 = $5,750 Favorable

Direct labor rate variance = Actual hours*Standard rate - Actual hours*Actual rate

Direct labor rate variance = 4,340*$40 - 4,340*$41.50

Direct labor rate variance = $173,600 - 180,110 = $6,510 Unfavorable

Direct labor efficiency variance = Standard hours*Standard rate - Actual hours*Standard rate

Direct labor efficiency variance = 920*5*$40 - 4,340*$40

Direct labor efficiency variance = $184,000 - 173,600 = $10,400 Favorable

ii. No, management did not make good decision as both the purchase price variance for aluminium and labor rate variance are unfavorable.